# On the jump of the cover time in random geometric graphs

### v2

@@ Abstract.
In this paper we study the cover time of the simple random walk on the giant component of supercritical $d$-dimensional random geometric graphs on $n$ vertices. We show that the cover time undergoes a jump at the connectivity threshold radius $r_{c}$: with $r_{g}$ denoting the threshold for having a giant component, we show that if the radius $r$ satisfies $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$ for $\varepsilon>0$ arbitrarily small, the cover time of the giant component is asymptotically almost surely $\Theta(n\log^{2}n$). On the other hand, we show that for $r\geq(1+\varepsilon)r_{c}$, the cover time of the graph is asymptotically almost surely $\Theta(n\log n)$ (which was known for $d=2$ only for a radius larger by a constant factor).
@@ 1. Introduction
The simple random walk on the vertices of a (throughout the paper finite and connected) graph $G=(V,E)$ is a discrete-time stochastic process with random variables $X_{0},X_{1},\ldots$ such that $X_{0}=v$ for some $v\in V$, and for $i\geq 0$, $X_{i+1}$ is a vertex chosen uniformly at random from the neighbors of $X_{i}$. The cover time of the simple random walk is then defined as
$\min\{t\geq 0:\bigcup_{s=0}^{t}\{X_{s}\}=V\}.$
The cover time in general depends on the starting vertex, and it therefore makes sense to consider
$\max_{v\in V,X_{0}=v}\min\{t\geq 0:\bigcup_{s=0}^{t}\{X_{s}\}=V\}.$
Also, since the cover time is a random variable, we are interested in its expected value (starting from a worst-case vertex). We thus define for $G(V,E)$,
$\tau_{cov}(G)=\mathbb{E}[\max_{v\in V,X_{0}=v}\min\{t\geq 0:\bigcup_{s=0}^{t}\{X_{s}\}=V\}].$
@@ 1.1. Main results
The cover time of $\mathcal{G}(n,r,d)$ was, as far as we know, only studied in two papers: first, Avin and Ercal (see [5]) showed that for $G\in\mathcal{G}(n,r,2)$ with $r^{2}>8\log n$, that is, a constant factor above the connectivity threshold, $\tau_{cov}(G)=\Theta(n\log n)$ a.a.s. Second, Cooper and Frieze [12] showed that for $d\geq 3$ and $G\in\mathcal{G}(n,r,d)$ with $r=(c\log n/V_{d})^{1/d}$, with $c>1$ being arbitrary and $V_{d}$ as before being the volume of the $d$-dimensional unit ball, a.a.s. $\tau_{cov}=c\log\left(\frac{c}{c-1}\right)n\log n$ (their result does not apply for $d=2$). In this paper, we look at sparser graphs not dealt with by these papers: for $G\in\mathcal{G}(n,r,d)$ and $r\geq r_{g}$, we denote by $L_{1}(G)$ its largest component (that is, the connected graph induced by the vertex set of this component). Our first main result is the following:
@@
Theorem 1.1 and 1.2 together with known results for $d\geq 3$ show that for $G\in\mathcal{G}(n,r,d)$, $\tau_{cov}$ undergoes a jump at $r_{c}$, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$. Our proofs shed some light into the behavior of $\tau_{cov}$ right at $r_{c}$, and we comment on this in the concluding remarks (we do not know the exact behavior, though).
@@ 1.2. Further related work
In the context of a percolated grid (of side length $n$), some results are known as well: whereas the cover time of (unpercolated) $\mathbb{Z}^{2}$ is a.a.s. of the order $n^{2}\log^{2}n$ and of (unpercolated) $\mathbb{Z}^{d}$ for $d\geq 3$ of the order $n^{d}\log n$, Abe in [1] showed that for any $d\geq 2$, the cover time of the giant component of supercritical $\mathbb{Z}^{d}$ is a.a.s. of the order $n^{2}\log^{2}n$, thereby showing a quantitative difference of unpercolated and percolated $\mathbb{Z}^{d}$ for $d\geq 3$.
@@ 1.3. Outline of proofs
The lower bound of the proof of Theorem 1.1 is comparably simpler: depending on the regime of the radius, either we find pending paths of order $\log n/r^{d}$ (as in the paper of Abe [1]), or we will find many vertices of degree $1$ inside the giant component. The upper bound is also inspired by Abe’s result in [1] on $\tau_{cov}$ of a supercritical percolation cluster on $\mathbb{Z}^{d}$: we first mimic Abe’s backbone construction from [1] over Bernoulli percolation now using a suitable renormalization scheme on the random geometric graph for fixed radius $r$. We then use this structure to build a flow that dissipates energy in a convenient way: we have to be careful with the construction as there are additional obstacles that do not appear in Bernoulli percolation, such as the fact that crossings could be arbitrarily close to one another or that we may have to extract induced paths from paths in densely populated areas. We then extend this result to larger radii below the connectivity threshold by a simple coupling. We remark that if we were interested in showing Theorem 1.1 only for $r=\Theta(1)$ and $r\geq Cr_{g}$ for $C$ being a large constant, the theorem would directly follow as a direct application of [1], using a tessellation of cubes of side length equal to $r/\sqrt{5}$ (for $d=2$, and changing the constant for other dimensions) and a coupling with Bernoulli percolation. However, since we show it for $r\geq(1+\varepsilon)r_{g}$, we need the already mentioned renormalization scheme for the upper bound.
The proof of Theorem 1.2 elaborates on the idea of the upper bound of Theorem 1.1 by additional ingredients: on the one hand, since the typical degree is high, one extra splitting of the flow is introduced in order to be able to bound the effective resistance between any two vertices of a certain backbone (which is large and well spread out) by $O(1/\log n)$. On the other hand, we then show that other vertices outside this backbone cannot be separated from the backbone by taking out too few vertices: we will find many short internally vertex-disjoint paths between any vertex outside and some vertex inside the backbone, on which we can split the flow again evenly.
@@
Plan of paper. In Section 2 we recall basic facts on random walks, in particular on the resistance, as well as facts on random geometric graphs and Bernoulli percolation. In Section 3 we then introduce a renormalization scheme that will turn out to be very useful in our proofs. Section 4 is then dedicated to the proof of the lower bound of Theorem 1.1. In Section 5 we then show the corresponding upper bound of Theorem 1.1, first for $r=\Theta(1)$, and later we show the small modifications needed for $r=\omega(1)$. Finally, in Section 6 we prove Theorem 1.2.
@@ 2.2. Random geometric graphs
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r>r_{g}$ and also $r=O(1)$. A.a.s. the following is true: there exist positive constants $C=C(r,d)$ and $C^{\prime}=C^{\prime}(r,d)$, such that for any two vertices $x,y\in L_{1}(G)$ with $d_{E}(x,y)\geq C\log n$, we have $d_{G}(x,y)\leq C^{\prime}d_{E}(x,y)$.
We get the following immediate corollary:
@@
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r>r_{g}$ and also $r=O(1)$. Let $V^{\prime}\subseteq L_{1}(G)$ be such that for any vertex $x\in L_{1}(G)$ we have that $d_{E}(x,V^{\prime})<C\log n$ for the same constant $C$ as in Theorem 2.11. Then, a.a.s., there exists a constant $c^{\prime}>0$ such that $d_{G}(x,V)\leq c^{\prime}\log n$ for any vertex $x\in L_{1}(G)$.
@@
On the one hand, for every $x\in L_{1}(G)$ there exists $y\in V^{\prime}$ with $d_{E}(x,y)<C\log n$, but on the other hand, since $x\in L_{1}(G)$, a.a.s. there exists $v\in L_{1}(G)$ with $d_{E}(x,v)\geq n^{\varepsilon}$ for some $\varepsilon>0$. Thus, there exists $z\in L_{1}(G)$ such that $d_{E}(x,z)\geq C\log n$ and $d_{E}(x,z)\leq 2C\log n$: indeed, consider a shortest path $x=v_{0},v_{1},\ldots,v_{k}=v$ for some $k\in\mathbb{N}$: since the Euclidean distance between two consecutive vertices is at most $r$, there exist vertices $v_{i}$ and $v_{i+1}$ on this path such that $d_{E}(x,v_{i})<C\log n$ and $d_{E}(x,v_{i+1})\geq C\log n$, but at the same time $d_{E}(x,v_{i+1})\leq C\log n+r\leq 2C\log n$. Therefore, by Theorem 2.11, $d_{G}(x,v_{i+1})\leq c^{\prime}\log n$ for some $c^{\prime}>0$ and since $v_{i+1}\in V^{\prime}$, the statement follows. ∎
@@ 2.3. Bernoulli percolation
We also need results from Bernoulli percolation. Consider an i.i.d. family of independent Bernoulli variables with mean $p$, indexed by the edges of $\mathbb{Z}^{d}$ (typically called bonds). We say that the bond $e$ is open if its associated variable is equal to 1. For a given realization, we generate a subgraph of $\mathbb{Z}^{d}$ with edges given by the open bonds, this random graph model is known as Bernoulli bond percolation. For $d\geq 2$, it is known that this model undergoes a non-trivial phase transition: there exists $10<p_{c}^{\mbox{bond}}<$ 1 such that if $p>p_{c}^{\mbox{bond}}$ then a.s. there is a unique (maximal) infinite connected component $\mathcal{C}$ (see for example [9]), which we call the infinite cluster. If one considers variables indexed by the vertices of $\mathbb{Z}^{d}$ instead of bonds we have Bernoulli site percolation. We say that a vertex is open if its associated variable is 1. We say that a path of vertices is open if all its vertices are open. As in bond percolation, for $d\geq 2$ there exists $10<p_{c}^{\mbox{site}}<$ 1 such that if $p>p_{c}^{\mbox{site}}$ then there is a.s. a unique infinite cluster.
We denote by $\mathbb{P}^{\mbox{site}}_{p}$ the law of Bernoulli site percolation with parameter $p$, and by $\mathbb{P}^{\mbox{bond}}_{p}$ the law of Bernoulli bond percolation with parameter $p$, respectively. We denote by $L_{1}^{\mbox{bond}}$ the largest connected component of Bernoulli bond percolation inside the box $\left[-n^{1/d}/2,n^{1/d}/2\right]^{d}$ if it exists. Equipping $L_{1}^{\mbox{bond}}$ with the network structure induced by the simple random walk, we have the following sharp estimate for the effective resistance of this network:
@@
Let $d\geq 2$. For bond percolation on $\mathbb{Z}^{d}$ and $p\in(p_{c}^{\mbox{bond}},1)$ there are constants $c_{1},c_{2}>0$ such that a.a.s., it is true that
$c_{1}\log n\leq\max_{x,y\in L_{1}^{\mbox{bond}}}\mathcal{R}\left(x\longleftrightarrow y,L_{1}^{\mbox{bond}}\right)\leq c_{2}\log n.$ (2.10)
@@
Consider site percolation on $\mathbb{Z}^{2}$ with parameter $p>p_{c}^{\mbox{site}}$. There exist positive constants $c_{1},c_{2},c_{3}$ such that
$\begin{split}\mathbb{P}^{\mbox{site}}_{p}\big{(}&\text{There exist at least $c_{1}n$ disjoint horizontal crossings}\\ &\text{ in $([0,m]\times[0,n])\cap\mathbb{Z}^{2}$}\big{)}\geq 1-c_{2}me^{-c_{3}n}.\end{split}$ (2.11)
@@
Let $d\geq 2$ and $p>p_{c}^{\mbox{site}}$. There exist constants $\kappa_{0},c$ such that a.a.s. the following holds: for any $x,y\in L_{1}^{\mbox{bond}}$, and any $\kappa>\kappa_{0}$, if $d_{E}(x,y)\leq\kappa\log n$, then $d_{G}(x,y)\leq c\kappa\log n.$
@@ 3. Renormalization scheme
Now, let $\mu=\mu(d)$ be largest constant so that any two vertices in adjacent cubes (that is, sharing a $(d-1)-$dimensional face) of side length $r/\mu$ are connected by an edge (for example, in $d=2$, $\mu=\sqrt{5}$). We will consider a tessellation using such cubes $\mathrm{Q_{i}}$ later on as well, for now we use it do define crossings. Let $U=[a_{1},a_{2}]\times[b_{1},b_{2}]\times U_{d-2}$ be a parallelepiped in $\mathbb{R}^{d}$, with $U_{d-2}$ a parallelepiped in $\mathbb{R}^{d-2}$. We say that the parallelepiped $U$ has a crossing component, if there exists a connected component (of the underlying $G\in\mathcal{G}(n,r,d)$) within the parallelepiped that contains vertices at Euclidean distance at most $\frac{r}{\mu}$ from each of the faces of the parallelepiped. Similarly as we did for $\mathbb{Z}^{d}$, we define also crossings: a horizontal crossing of $G\in\mathcal{G}(n,r,d)$ in $U$ is a self-avoiding path of $G$ such that the vertices of the path are totally contained in $U$, the first vertex of the path has first coordinate in the interval $(a_{1},a_{1}+r/\mu)$, and the last vertex of the path has first coordinate in the interval $(a_{2}-r/\mu,a_{2})$. A vertical crossing is defined in the same way, exchanging the first coordinate by the second coordinate and $a_{1},a_{2}$ by $b_{1},b_{2}$, and as before, we use also the definition of a long crossing for a crossing in the longer dimension. We remark that while a horizontal (or vertical) crossing is a self-avoiding path, a crossing component is in general not a path (but it contains such a path).
@@
Let $d\geq 2$, let $M$ be a large enough constant, let $G\in\mathcal{G}(n,r,d)$ and let $r\geq(1+\varepsilon)r_{g}$ for arbitrarily small $\varepsilon>0$. Fix $\mathrm{i}\sim\mathrm{j}\in\mathbb{Z}^{d}$. There exists a constant $\gamma>0$ such that
@@
A 2-dimensional slice of $\mathcal{K}_{n}$ is a subset of cubes with indices $\mathrm{i}$ such that all but $2$ coordinates are fixed (see Figure 2 for an example). We denote such a slice by $\mathcal{K}_{n}^{2}$. Since very $2$-dimensional slice is isomorphic to $\mathbb{Z}^{2}$ restricted to a box of a certain side length, by abuse of notation we can thus also define both vertical and horizontal orientations for each slice. We remark also, that, as before, the good cubes of $\mathcal{K}^{2}_{n}$ dominate supercritical 2-dimensional site percolation restricted to a large box.
@@
Let $d\geq 2$ and $\varepsilon>0$, let $G\in\mathcal{G}(n,r,d)$ and assume $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$ and assume $M$ large enough. Let $\alpha$ be a large enough positive constant. Then, a.a.s. there exists $c>0$ such that each $\alpha$-logarithmic strip that is not in the boundary of $\Lambda_{n}$ of every 2-dimensional slice of $\Lambda_{n}$ has at least $c\log n$ disjoint long crossings of good cubes.

### v3

@@ Abstract.
In this paper we study the cover time of the simple random walk on the giant component of supercritical $d$-dimensional random geometric graphs on $\mathrm{Poi}(n)$ vertices. We show that the cover time undergoes a jump at the connectivity threshold radius $r_{c}$: with $r_{g}$ denoting the threshold for having a giant component, we show that if the radius $r$ satisfies $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$ for $\varepsilon>0$ arbitrarily small, the cover time of the giant component is asymptotically almost surely $\Theta(n\log^{2}n$). On the other hand, we show that for $r\geq(1+\varepsilon)r_{c}$, the cover time of the graph is asymptotically almost surely $\Theta(n\log n)$ (which was known for $d=2$ only for a radius larger by a constant factor). Our proofs also shed some light onto the behavior around $r_{c}$.
@@ 1. Introduction
The simple random walk on the vertices of a (throughout the paper finite and connected) graph $G=(V,E)$ is a discrete-time stochastic process with random variables $X_{0},X_{1},\ldots$ such that $X_{0}=v$ for some $v\in V$, and for $i\geq 0$, $X_{i+1}$ is a vertex chosen uniformly at random from the neighbors of $X_{i}$. For a graph $G(V,E)$, the cover time of the simple random walk is then defined as
$\tau_{cov}(G)=\max_{v\in V,X_{0}=v}\mathbb{E}[\min\{t\geq 0:\bigcup_{s=0}^{t}\{X_{s}\}=V\}].$
In words, the cover time is the expected time it takes to visit all vertices, when starting from a worst-case vertex.
@@ 1.1. Main results
The cover time of $\mathcal{G}(n,r,d)$ was, as far as we know, only studied in two papers: first, Avin and Ercal (see [5]) showed that for $G\in\mathcal{G}(n,r,2)$ with $r^{2}>8\log n$, that is, a constant factor above the connectivity threshold, $\tau_{cov}(G)=\Theta(n\log n)$ a.a.s. Second, Cooper and Frieze [12] showed that for $d\geq 3$ and $G\in\mathcal{G}(n,r,d)$ with $r=(c\log n/V_{d})^{1/d}$, with $c>1$ being arbitrary and $V_{d}$ as before being the volume of the $d$-dimensional unit ball, a.a.s. $\tau_{cov}=c\log\left(\frac{c}{c-1}\right)n\log n$ (their result does not apply for $d=2$). In this paper, we look at sparser graphs not dealt with by these papers: for $G\in\mathcal{G}(n,r,d)$ and $r\geq r_{g}$, we denote by $L_{1}(G)$ its largest component (that is, the largest connected graph induced by the vertex set of this component). Our first main result is the following:
@@
Theorem 1.1 and 1.2, together with known results for $d\geq 3$, show that for $G\in\mathcal{G}(n,r,d)$, $\tau_{cov}$ undergoes a jump at $r_{c}$, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$. Our proofs shed some light onto the behavior of $\tau_{cov}$ right at $r_{c}$, and we comment on this in the concluding remarks (we do not know the exact behavior, though).
@@ 1.2. Further related work
In the context of a percolated grid (of side length $n$), some results are known as well: whereas the cover time of (unpercolated) $\mathbb{Z}^{2}$ is a.a.s. of the order $n^{2}\log^{2}n$ and of (unpercolated) $\mathbb{Z}^{d}$ for $d\geq 3$ of the order $n^{d}\log n$, Abe in [1] showed that for any $d\geq 2$, the cover time of the giant component of supercritical Bernoulli percolation in $\mathbb{Z}^{d}$ is a.a.s. of the order $n^{2}\log^{2}n$, thereby showing a quantitative difference of unpercolated and percolated $\mathbb{Z}^{d}$ for $d\geq 3$.
@@ 1.3. Outline of proofs
The approach we follow is based on the idea of seeing the graph as an electric network. The main focus of the paper is therefore to obtain good enough bounds on the effective resistance.
The lower bound of the proof of Theorem 1.1 is comparably simpler: depending on the regime of the radius, either we find pending paths of order $\log n/r^{d}$ (as in the paper of Abe [1]), or we will find many vertices of degree $1$ inside the giant component. The upper bound is also inspired by Abe’s result in [1] on $\tau_{cov}$ of a supercritical percolation cluster on $\mathbb{Z}^{d}$: we first mimic Abe’s backbone construction from [1] over Bernoulli percolation now using a suitable renormalization scheme on the random geometric graph for fixed radius $r$. We then use this structure to build a flow that dissipates energy in a convenient way: we have to be careful with the construction as there are additional obstacles that do not appear in Bernoulli percolation, such as the fact that crossings could be arbitrarily close to one another or that we may have to extract induced paths from paths in densely populated areas. When the typical degree is tending to infinity, one extra splitting of the flow is introduced in order to bound the effective resistance between any two vertices of a certain backbone (which is large and well spread out) by $O(\log n/r^{2d})$.
The proof of Theorem 1.2 elaborates on the idea of the upper bound of Theorem 1.1 by an additional ingredient: on the one hand, since the typical degree is high, we need again the extra splitting from the case when the radius tends to infinity. On the other hand, we then show that other vertices outside this backbone cannot be separated from the backbone by taking out too few vertices: we will find many short internally vertex-disjoint paths between any vertex outside and some vertex inside the backbone, on which we can split the flow again evenly.
@@
Plan of paper. In Section 2 we recall basic facts on random walks, in particular on the effective resistance, as well as facts on random geometric graphs and Bernoulli percolation. In Section 3 we then introduce a renormalization scheme that will turn out to be very useful in our proofs. Section 4 is then dedicated to the proof of the lower bound of Theorem 1.1. In Section 5 we then show the corresponding upper bound of Theorem 1.1 for $r=\Theta(1)$. In Section 6 we extend this to the case $r=\omega(1)$. Finally, in Section 7 we prove Theorem 1.2.
@@ 2.2. Random geometric graphs
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r_{g}<r$. A.a.s. the following is true: there exist positive constants $C=C(r,d)$ and $C^{\prime}=C^{\prime}(r,d)$, such that for any two vertices $x,y\in L_{1}(G)$ with $d_{E}(x,y)\geq C\frac{\log n}{r^{d-1}}$, we have $d_{G}(x,y)\leq C^{\prime}d_{E}(x,y)/r$.
We get the following corollary:
@@
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r_{g}<r=O(r_{c})$. Let $V^{\prime}\subseteq L_{1}(G)$ be such that for any vertex $x\in L_{1}(G)$ we have that $d_{E}(x,V^{\prime})<C\log n/r^{d-1}$ for the same constant $C$ as in Theorem 2.11. Then, a.a.s., there exists a constant $c^{\prime}>0$ such that $d_{G}(x,V^{\prime})\leq c^{\prime}\log n/r^{d}$ for any vertex $x\in L_{1}(G)$.
@@
On the one hand, for every $x\in L_{1}(G)$ there exists $y\in V^{\prime}$ with $d_{E}(x,y)<C\log n/r^{d-1}$, but on the other hand, since $x\in L_{1}(G)$, a.a.s. there exists $v\in L_{1}(G)$ with $d_{E}(x,v)\geq n^{\varepsilon}$ for some $\varepsilon>0$. Thus, there exists $z\in L_{1}(G)$ such that $d_{E}(x,z)\geq C\log n/r^{d-1}$ and $d_{E}(x,z)=O(\log n/r^{d-1})$: indeed, consider a shortest path $x=v_{0},v_{1},\ldots,v_{k}=v$ for some $k\in\mathbb{N}$: since the Euclidean distance between two consecutive vertices is at most $r$, there exist vertices $v_{i}$ and $v_{i+1}$ on this path such that $d_{E}(x,v_{i})<C\log n/r^{d-1}$ and $d_{E}(x,v_{i+1})\geq C\log n/r^{d-1}$, but at the same time $d_{E}(x,v_{i+1})\leq C\log n/r^{d-1}+r=O(C\log n/r^{d-1})$, where the equality holds from our assumption $r=O(r_{c})$. Therefore, by Theorem 2.11, $d_{G}(x,v_{i+1})\leq c^{\prime}\log n/r^{d}$ for some $c^{\prime}>0$ and since $v_{i+1}\in V^{\prime}$, the statement follows. ∎
@@ 2.3. Bernoulli percolation
We also need results from Bernoulli percolation. Consider an i.i.d. family of independent Bernoulli variables with mean $p$, indexed by the edges of $\mathbb{Z}^{d}$ (typically called bonds). We say that the bond $e$ is open if its associated variable is equal to 1. For a given realization, we generate a subgraph of $\mathbb{Z}^{d}$ with edges given by the open bonds, this random graph model is known as Bernoulli bond percolation. For $d\geq 2$, it is known that this model undergoes a non-trivial phase transition: there exists $10<p_{c}^{bond}<$ 1 such that if $p>p_{c}^{bond}$ then a.s. there is a unique (maximal) infinite connected component $\mathcal{C}$ (see for example [9]), which we call the infinite cluster. If one considers variables indexed by the vertices of $\mathbb{Z}^{d}$ instead of bonds we have Bernoulli site percolation. We say that a vertex is open if its associated variable is 1. We say that a path of vertices is open if all its vertices are open. As in bond percolation, for $d\geq 2$ there exists $10<p_{c}^{site}<$ 1 such that if $p>p_{c}^{site}$ then there is a.s. a unique infinite cluster.
We denote by $\mathbb{P}^{site}_{p}$ the law of Bernoulli site percolation with parameter $p$, and by $\mathbb{P}^{bond}_{p}$ the law of Bernoulli bond percolation with parameter $p$, respectively. We denote by $\mathcal{C}^{bond}(n)$ the largest connected component of Bernoulli bond percolation inside the box $\left[-n^{1/d}/2,n^{1/d}/2\right]^{d}$ if it exists. Equipping $\mathcal{C}^{bond}(n)$ with the network structure induced by the simple random walk, we have the following sharp estimate for the effective resistance of this network:
@@
Let $d\geq 2$. For bond percolation on $\mathbb{Z}^{d}$ and $p\in(p_{c}^{bond},1)$ there are constants $c_{1},c_{2}>0$ such that a.a.s., it is true that
$c_{1}\log n\leq\max_{x,y\in L_{1}^{bond}}\mathcal{R}\left(x\longleftrightarrow y,\mathcal{C}^{bond}(n)\right)\leq c_{2}\log n.$ (2.10)
@@
Consider site percolation on $\mathbb{Z}^{2}$ with parameter $p>p_{c}^{site}$. There exist positive constants $c_{1},c_{2},c_{3}$ such that
$\begin{split}\mathbb{P}^{site}_{p}\big{(}&\text{There exist at least $c_{1}n$ disjoint horizontal crossings}\\ &\text{ in $([0,m]\times[0,n])\cap\mathbb{Z}^{2}$}\big{)}\geq 1-c_{2}me^{-c_{3}n}.\end{split}$ (2.11)
@@
Let $d\geq 2$ and $p>p_{c}^{site}$. There exist constants $\kappa_{0},c$ such that a.a.s. the following holds: for any $x,y\in\mathcal{C}^{site}(n)$, and any $\kappa>\kappa_{0}$, if $d_{E}(x,y)\leq\kappa\log n$, then $d_{G}(x,y)\leq c\kappa\log n.$
@@ 3. Renormalization scheme
Now, let $\mu=\mu(d)$ be the largest constant so that any two vertices in adjacent cubes (that is, sharing a $(d-1)-$dimensional face) of side length $r/\mu$ are connected by an edge (for example, in $d=2$, $\mu=\sqrt{5}$). We will consider a tessellation using such cubes $\mathrm{Q_{i}}$ of side length $r/\mu$ later on as well, for now we use it do define crossings. Let $U=[a_{1},a_{2}]\times[b_{1},b_{2}]\times U_{d-2}$ be a parallelepiped in $\mathbb{R}^{d}$, with $U_{d-2}$ a parallelepiped in $\mathbb{R}^{d-2}$. We say that the parallelepiped $U$ has a crossing component, if there exists a connected component (of the underlying $G\in\mathcal{G}(n,r,d)$) within the parallelepiped that contains vertices at Euclidean distance at most $\frac{r}{\mu}$ from each of the faces of the parallelepiped. Similarly as we did for $\mathbb{Z}^{d}$, we define also crossings: a horizontal crossing of $G\in\mathcal{G}(n,r,d)$ in $U$ is a self-avoiding path of $G$ such that the vertices of the path are totally contained in $U$, the first vertex of the path has first coordinate in the interval $(a_{1},a_{1}+r/\mu)$, and the last vertex of the path has first coordinate in the interval $(a_{2}-r/\mu,a_{2})$. A vertical crossing is defined in the same way, exchanging the first coordinate by the second coordinate and $a_{1},a_{2}$ by $b_{1},b_{2}$, and as before, we use also the definition of a long crossing for a crossing in the longer dimension. We remark that while a horizontal (or vertical) crossing is a self-avoiding path, a crossing component is in general not a path (but it contains such a path).
@@
Let $d\geq 2$, let $M$ be a large enough constant, let $G\in\mathcal{G}(n,r,d)$ and let $r\geq(1+\varepsilon)r_{g}$ for arbitrarily small $\varepsilon>0$. Fix $\mathrm{i}\in\mathbb{Z}^{d}$. If $\mathrm{i}\sim\mathrm{j}$, there exists a constant $\gamma>0$ such that
@@
A 2-dimensional slice of $\mathcal{K}_{n}$ is a subset of cubes with labels $\mathrm{i}$ such that all but $2$ coordinates are fixed (see Figure 2 for an example). We denote such a slice by $\mathcal{K}_{n}^{2}$. Since very $2$-dimensional slice is isomorphic to $\mathbb{Z}^{2}$ restricted to a box of a certain side length, by abuse of notation we can thus also define both vertical and horizontal orientations for each slice. We remark also, that, as before, the good cubes of $\mathcal{K}^{2}_{n}$ dominate supercritical 2-dimensional site percolation restricted to a large box.
@@
Let $d\geq 2$ and $\varepsilon>0$, let $G\in\mathcal{G}(n,r,d)$ and assume $(1+\varepsilon)r_{g}\leq r=O(r_{c})$ and assume $M$ large enough. Let $\alpha$ be a large enough positive constant. Then, a.a.s. there exists $c>0$ such that each $\alpha$-logarithmic strip that is not in the boundary of $\Lambda_{n}$ of every 2-dimensional slice of $\Lambda_{n}$ has at least $c\log n$ disjoint long crossings of good cubes.